OpenStudy (anonymous):

what is the area of an equilateral triangle with radius 15

OpenStudy (anonymous):

do you mean equialateral triangle side = 15

OpenStudy (anonymous):

if it is side and equals to 15, formula for area of equilateral triangle = (side 1 )(side2)/2 = 15*15/2 = 112.5 unit square. so area of equilateral triangle is 112.5 unit square.